OpenStudy (jilltalks):

Pyramid A is a square pyramid with a base side length of 12 inches and a height of 8 inches. Pyramid B is a square pyramid with a base side length of 24 inches and a height of 16 inches. How many times bigger is the volume of pyramid B than pyramid A?

OpenStudy (wolf1728):

The volume of a pyramid is (length * width * height) / 3 So for pyramid A volume = (12 * 12 * 8) / 3 For pyramid B volume = (24 * 24 * 16) / 3 That should get you off to a good start.

OpenStudy (jilltalks):

8?

OpenStudy (jilltalks):

it was 8 thank you